# Page Pages

### v1

@@ 1 Introduction
Two- and three-point correlation functions ($2/3$PCF) are widely used in cosmology. Algorithms to compute these correlation functions first cluster their objects into representative points to ease the computational load of computing $\mathcal{O}(n^{2})$ pairwise distances. To numerically compute the correlation function, the pairwise distance between all of the points is then computed, and an estimator is computed on all of the points in each distance bin.
@@
A long-standing goal in the computational sciences is to reconcile epistemic (model) uncertainties from aleatoric (data) uncertainties (Osband et al., 2023; Hüllermeier & Waegeman, 2021). Aleatoric uncertainties describe uncertainty that is inherent to a statistic. This can be due to data quality limitations or inherent stochasticity of the quantity one is trying to model. On the other hand, epistemic uncertainties reflect the inability to generalize from a finite data set due to inherent limitations of the model. In the case of $2$PCFs and $3$PCFs, the epistemic uncertainty is often overlooked: the best assignment of an object to a given cluster is often unclear, as data points that are “in-between” two or more learned cluster centers can reasonably be chosen to fall in any of the clusters. This ambiguity introduces an epistemic uncertainty, as the model will struggle to appropriately assign objects to representative clusters without enough data. Prior works usually assume this clustering uncertainty to be trivial, arguing that for a significantly large sample, the number of overestimated and underestimated distances are roughly equal, and so the errors tend to cancel out (Jarvis et al., 2004). For example, this was assumed for the Dark Energy Survey (DES) point spread function (PSF) modeling efforts (Jarvis et al., 2021). A PSF is an impulse response of an optical system to light, and the quality of a PSF model is often assessed with $\rho$ statistics (Rowe, 2010; Vogelsberger et al., 2016), which describe correlations in size and shape residuals of a PSF model relative to observed stars. These empirical PSF models are often trained using point source catalogs that are generated by SourceExtractor++ (Bertin & Arnouts, 1996) and filtered to be within a given size/magnitude range. DES covered a vast survey area and included millions of objects which were used to be fit and evaluate PSF models (Dark Energy Survey Collaboration, 2024; Jarvis et al., 2021). With so much data at hand, it was fair to assume that the number of overestimated and underestimated distances were roughly equal, and that downstream modeling errors from clustering were negligible. Surveys such as DES therefore look exclusively at uncertainties quantified via bootstrapping or jackknife (Jarvis et al., 2021; Jarvis & collaborators, 2024), which captures only the model’s robustness to individual data points. This is closer in nature to being an aleatoric uncertainty, and we will refer to it as aleatoric uncertainty throughout.
@@
This work presents a method to quantify epistemic uncertainty in regimes where one cannot make the same assumptions as Jarvis et al. (2004) due to the small number of data points available. For instance, clustering errors cannot be assumed to be negligible for the NIRCam PSF modeling efforts with COSMOS-Web (Casey et al., 2023), where the number of point sources is on the order of a few $100$ across the entire field of view (Berman et al., 2024). Even with distance offsets mitigated to a single bin with the TreeCorr algorithm (Jarvis et al., 2004), results can still be heavily biased. A toy example of this with the position-position autocorrelation is presented in Figure 2. The unbiased estimator for this correlation is the Landy–Szalay estimator (Landy & Szalay, 1993) shown in Equation 6.
@@
Our paper executes three experiments that are designed to explore each of the three themes outlined in our title: model uncertainty, differentiability, and surrogates.
@@
    Model Uncertainty and PSFs (§$\S$5): For this experiment, we use ShOpt.jl (Berman et al., 2024; Berman & McCleary, 2024) to produce PSF models for the COSMOS-Web NIRCam images (Casey et al., 2023). Using our proposed algorithm, we then compute $\rho$ statistics. Finally, we assess the epistemic uncertainty associated with our model and compare it to the uncertainty we would get with traditional bootstrapping techniques. Our PSF modeling experiment is geared towards observers, who do not have access to differentiable forward models of PSF residuals.
@@
    Differentiability and Gravity Simulations (§$\S$6): For this experiment, we forward model gravitational interactions using a series of ordinary differential equations. We then compute correlation functions on the resulting catalog and test different techniques for differentiating the correlation function outputs with respect to the cosmological parameters that determined the simulation. Our gravitational simulation experiment is designed with theorists in mind. Theorists can produce multiple realizations of their model through simulations that are derived from different cosmological parameters. Here, we assess the feasibility of differentiating a correlation function with respect to $G$.
@@
    Surrogates and Galaxy Intrinsic Alignment (§$\S$7): For this experiment, we use an existing surrogate model to relate astrophysical model parameters to IA (Pandya et al., 2024, 2025). Using this surrogate, we use Hamiltonian Monte Carlo (HMC) to find posterior distributions over astrophysical model parameters that are likely to correspond to correlations of zero. Our IA experiment shows how to compensate for forward models and estimators that are not easily differentiated, and are again geared towards theorists who are able to realize many instances of their forward model through inference in order to train a surrogate. Our usage of HMC leverages the differentiability of neural networks to accelerate the sampling procedure in comparison to traditional algorithms like Metropolis–Hastings (Betancourt, 2017; Robert et al., 2004).
The remainder of the paper is organized as follows: §$\S$2 discusses the relevant notation and theory for $\rho$ statistics and IA, §$\S$3 follows with an outline of the data products used in our study, §$\S$4 lays out additional notation as well as details for the algorithm used in this paper, and §$\S$5 -7 presents the results of the experiments described above followed by our conclusions in §$\S$8.
@@ 2.1 PSF Modeling
where $e_{\rm PSF}$ is the ellipticity of the real PSF, i.e., the star ellipticity, $T_{\rm PSF}$ is the size of the real PSF, $\delta e_{\rm PSF}$ is the difference between the ellipticity of the real and model PSFs at position $\bm{x}$, and $\delta T_{\rm PSF}$ is the difference between the sizes of the real and model PSFs at position $\bm{x}$. Brackets denote averages over all pairs within a separation $\bm{\theta}$, and asterisks denote complex conjugates.
@@
where $w_{i/j}$ represents the weight uncertainties of object $(i/j)$, $\epsilon_{t,i/j}$ represents tangential shear of object $(i/j)$ and $\epsilon_{\times,i/j}$ represents the cross shear of object $(i/j)$ (Kilbinger, 2015; Schneider et al., 2002). The complex components of the products in Equations 7 - 11 are known to average out to zero due to parity symmetry (Rowe, 2010; Schneider et al., 2006), and are thus ignored in the estimator.
@@ 2.2 Galaxy Intrinsic Alignment
Intrinsic alignment describe correlations between galaxy orientations and the underlying distribution of dark matter where they are embedded; see Lamman et al. (2023) for a review. The three relevant IA correlation functions are the galaxy position-position $\xi(r)$, position-orientation $\omega(r)$, and orientation-orientation $\eta(r)$ correlation functions, and are estimated according to Equations 14-16.
@@ 3.1 Point Source Catalog
The point sources used for generating our PSF models come from the SE+++++ + (Bertin & Arnouts, 1996) COSMOS-Web catalog (Casey et al., 2023, Shuntov et al. in prep). We restrict ourselves to imaging from the F115115115115W NIRCam (Rieke et al., 2003, 2005; Beichman et al., 2012) wavelength filter, as the PSFs in this wavelength are the closest to being well-approximated by a Gaussian (Berman et al., 2024). $\rho$ statistics and the related Hirata–Seljak–Mandelbaum (HSM) adaptive moments (Hirata & Seljak, 2003; Mandelbaum et al., 2005) are based off of the Gaussian approximation of the PSF, and so the use of $\rho$ statistics in our experiments is most appropriate in this wavelength. PSF models are obtained using ShOpt.jl (Berman et al., 2024; Berman & McCleary, 2024). Our choice of PSF fitter is discussed in more detail in Appendix A. Our ShOpt fits can be seen in Figure 3. The ellipticites of the objects in our data set are also visualized in Figure 4. We use the same configuration file as found in the appendix of Berman et al. (2024). Our data set statistics for PSF modeling are summarized in Table 1.
@@ 3.2 Simulated Galaxy Catalog
There are plenty of widely-used physically realistic N-body simulations, using publically available codes such as Gadget-2/3 (Springel, 2005), ART (Kravtsov et al., 1997), RAMSES (Teyssier, 2002), and Enzo (Brummel-Smith et al., 2019; Bryan et al., 2014), see Vogelsberger et al. (2020) for a review. However, this work aims only to have a simple differentiable forward model. Instead of using Gadget-2 or similar tools, we write our simulations from scratch in Julia and Python and time evolve objects according to ordinary differential equations from Newtonian mechanics. Our main implementation uses Jax (Bradbury et al., 2018) with Diffrax (Kidger, 2021). We specify starting positions and velocities for objects $[\vec{x},\vec{v}]$, which are then evolved via $\frac{d\vec{x}}{dt}=\vec{v}$, $\frac{d\vec{v}}{dt}=\vec{a}$, and $\vec{a}=-\frac{GM}{|\vec{x}|^{3}}\vec{x}$. We begin by randomly initializing galaxies with some mass $M$, velocity $\vec{v}$, and initial position $\vec{r}$. These galaxies exist in a gravitational field where the galaxy-galaxy gravitational attraction is assumed to be negligible compared to the attraction toward the center of the field. Thus, we can solve each ODE system in Diffrax independently and use the vmap function to efficiently time evolve each object. While our toy simulation is not very physically meaningful, this approach outputs catalogs that are differentiable with respect to the input cosmology, which in this case is just the gravitational constant $G$. The number of galaxies in the simulation can vary, allowing us to study the efficacy of our algorithm as a function of catalog sizes.
@@ 3.3 Intrinsic Alignment Catalog
[…]
The contributions of Van Alfen et al. (2024) ensures that the model takes into account IA, including the central alignment strength $\mu_{\rm cen}$ and the satellite alignment strength $\mu_{\rm sat}$.
For each of the seven input parameters, ten realizations of the model were produced. Once the catalogs were generated, the three IA correlation functions were estimated according to Equations 14 - 16. This led to a data matrix of the shape $(110,526;7;100;3;20)$. In other words, we have $110,526$ tuples of seven input values that produced ten realizations of three correlation functions that each occupy twenty radial bins. This catalog is used to train the IA emulator (IAEmu) surrogate model introduced in Pandya et al. (2024). IAEmu is a neural network based emulator designed to predict galaxy IA statistics from a given HOD, with the alignments parameterized according to the two parameter family $\mu_{cen}$ and $\mu_{sat}$ (Van Alfen et al., 2024). These $110,526$ tuples are broken down in training, validation, and testing sets for IAEmu. This is summarized in Table 2.
@@ 4 Algorithm Overview
In this section, we outline our algorithm used to estimate correlations. Our algorithm has two variants, one geared toward estimating clustering uncertainties and the other towards differentiability. Two variants are needed because estimating cluster uncertainty requires sampling, which is incompatible with differentiation. There are a variety of ways to adapt sampling to be a differentiable process, which we explore in depth in §6. Throughout this work, we use most of the formalism laid out in Bezdek et al. (1984). Our formalism differs in that it uses a distance function to compute angular distances on the sky rather than using a matrix to define a norm. We also define the quantity matrix, which allows us incorporate the quantities we wish to correlate into the original formalism. For details on the fuzzy sets formalism, see Zadeh (1965) and (page 38, Aluffi, 2021). Algorithm 1 first uses K-means+++++ + (Kapoor & Singhal, 2017) to initialize a series of cluster centers. This algorithm adds initial cluster centers sequentially by considering the distances to centers that have already been determined. This initialization has been shown to help clustering algorithms and is better than choosing initial cluster centers at random (Kapoor & Singhal, 2017). We next use the fuzzy-c-means algorithm (Bezdek et al., 1984) to obtain final cluster centers and weight assignments that minimize the functional
@@
While this does not have a compact matrix representation, it is stable at all numerical scales (Vincenty, 1975). In this way, we are sacrificing speed for stability. This design choice is discussed more in §$\S$6. The two argument $\arctan$ function is used to constrain the solution to the correct quadrant. Our usage of the Vincenty formula also assures that our distance estimates are consistent with Astropy (Robitaille et al., 2013), which also uses the Vincenty formula for this task.
@@
    Initialize centers via K-means+++++ + and weights via random initialization.
@@
1:  Initialize centers $v$ using K-Means+++++ +
@@
Figure 7: An example weighting for estimates for a distance bin bounded below by $1$ and above by $2$ for various sharpnesses.
@@
1:  Initialize centers $v$ using K-Means+++++ +
@@
5 Model Uncertainty
In this section, we discuss the results of Algorithm 1 on our dataset, which we compare with a bootstrapping baseline to obtain epistemic and aleatoric uncertainties.
6 Differentiability
For both approaches b and c (as defined in §$\S$4), the naive approach to computing the gradient $\frac{\partial\xi}{\partial\alpha_{i}}$ or $\frac{\partial\xi}{\partial\delta_{i}}$ would involve tracking the computation graph through the entire fuzzy-c-means algorithm, which can sometimes take over $100$ iterations. Another complication is that the distance calculations and normalization computations cannot be expressed through matrix multiplication, making it difficult to accelerate the gradient calculations with a GPU. For approach $b$, taking the gradients in this manner proved to be prohibitively expensive via reverse mode automatic differentiation, for the reasons outlined above. With forward mode differentiation, the gradient calculation was on the order of a minute. This is expected, since we are differentiating several output bins with respect to $1$ input (gravity).
7 Surrogates
In this section, we utilize the surrogate model developed in Pandya et al. (2024, 2025), IAEmu, which estimates galaxy IA correlations from an underlying set of HOD parameters. We use IAEmu to optimize model parameters that minimize the IA correlations, to study regions of parameter space that correspond to minimal IA contamination and are thus of physical interest. IAEmu serves as an emulator for Halotools-IA, which extends traditional HOD modeling to include IA information through a two parameter family, $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, which govern central and satellite alignment strengths, respectively. More specifically, IAEmu outputs the galaxy position-position ($\xi$), position-orientation ($\omega$), and orientation-orientation ($\eta$) correlations as well as estimates of the aleatoric uncertainty (shape noise) of each correlation. This noise is not physical, and corresponds to variance across realizations of the underlying HOD due to limited cosmological volumes when conducting the HOD simulations.
We pose this optimization problem because IA often act as a contaminant for weak lensing measurements, and moreover, there are known regions of parameter space that lead to low correlations. Thus, with this experiment, we are able to demonstrate that a differentiable forward model allows one to learn something about the underlying physics of IA. In other words, we are able to recover a known parameter space even without hard-coding the physical laws into the surrogate directly.  [^3] The exact correlation that you would subtract off is the gravitationally lensed shape - intrinsic shape $(GI)$ correlation (Lamman et al., 2023); however, the strength of that correlation is closely related to that of $\omega$.
To isolate regions of parameter space in $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ that minimize the IA correlation $\omega(r)$, we utilize Hamiltonian Monte Carlo (HMC). HMC is a gradient based Monte Carlo technique which leverages the differentiability of NNs and uses gradient information to efficiently traverse the input parameter space towards regions that maximize the posterior probability. For a general review of HMC, see Betancourt (2017). Our HMC experiment mirrors Pandya et al. 2025 (in prep). We run our HMC with a No U-Turn Sampler (NUTS) (Hoffman et al., 2014) with $1000$ warmup steps and $1000$ samples. We fix the HOD parameters according to the fiducial values present in IllustrisTNG300-1 (Nelson et al., 2019) and Table C1 of Van Alfen et al. (2024). This gives us three sets of HOD parameters, which each correspond to a different stellar mass cutoff in galaxies. We impose uniform priors of $[-1,1]$ on $\mu_{\rm cen}$ and $\mu_{\rm sat}$ and use a multivariate normal distribution centered at zero as our likelihood, with a covariance given by the IAEmu aleatoric uncertainty prediction. In this way, we encourage HMC to arrive at regions of parameter space that minimize the $\omega$ correlation, while also incorporating the inherent shape noise into the likelihood. A general outline of our procedure is given in algorithm 3.

### v3

@@ 1 Introduction
Two- and three-point correlation functions (2/32/3PCF) are widely used in cosmology. Algorithms to compute these correlation functions first cluster their objects into representative points to ease the computational load of computing $\mathcal{O}(n^{2})$ pairwise distances. To numerically compute the correlation function, the pairwise distance between all of the points is then computed, and an estimator is computed on all of the points in each distance bin.
@@
A long-standing goal in the computational sciences is to reconcile epistemic (model) uncertainties from aleatoric (data) uncertainties (Hüllermeier & Waegeman, 2021; Osband et al., 2023). Aleatoric uncertainties describe uncertainty that is inherent to a statistic. This can be due to data quality limitations or inherent stochasticity of the quantity one is trying to model. On the other hand, epistemic uncertainties reflect the inability to generalize from a finite data set due to inherent limitations of the model. In the case of 22PCFs and 33PCFs, the epistemic uncertainty is often overlooked: the best assignment of an object to a given cluster is often unclear, as data points that are “in-between” two or more learned cluster centers can reasonably be chosen to fall in any of the clusters. This ambiguity introduces an epistemic uncertainty, as the model will struggle to appropriately assign objects to representative clusters without enough data. With more data points, it is easier to identify the most appropriate way to cluster the objects and thus epistemic uncertainty should decrease in the main, though individual data points will still face ambiguity. Prior works usually assume this clustering uncertainty to be trivial, arguing that for a significantly large sample, the number of overestimated and underestimated distances are roughly equal, and so the errors tend to cancel out (Jarvis et al., 2004). For example, this was assumed for the Dark Energy Survey (DES) point spread function (PSF) modeling efforts (Jarvis et al., 2021). A PSF is an impulse response of an optical system to light, and the quality of a PSF model is often assessed with $\rho$ statistics (Rowe, 2010; Vogelsberger et al., 2016), which describe correlations in size and shape residuals of a PSF model relative to observed stars. These empirical PSF models are often trained using point source catalogs that are generated by SourceExtractor++ (Bertin & Arnouts, 1996) and filtered to be within a given size/magnitude range. DES covered a vast survey area and included millions of objects which were used to be fit and evaluate PSF models (Dark Energy Survey Collaboration, 2024; Jarvis et al., 2021). With so much data at hand, it was fair to assume that the number of overestimated and underestimated distances were roughly equal, and that downstream modeling errors from clustering were negligible. Surveys such as DES therefore look exclusively at uncertainties quantified via bootstrapping or jackknife (Jarvis et al., 2021; Jarvis & collaborators, 2024), which captures only the model’s robustness to individual data points. This is closer in nature to being an aleatoric uncertainty, and we will refer to it as aleatoric uncertainty throughout.
@@
This work presents a method to quantify epistemic uncertainty in regimes where one cannot make the same assumptions as Jarvis et al. (2004) due to the small number of data points available. For instance, clustering errors cannot be assumed to be negligible for the NIRCam PSF modeling efforts with COSMOS-Web (Casey et al., 2023), where the number of point sources is on the order of a few $100$ across the entire field of view (Berman et al., 2024). Even through the TreeCorr algorithm (Jarvis et al., 2004) makes sure distance offsets are less than a bin width, results can still be heavily biased. A toy example of this with the position-position autocorrelation is presented in Figure 2. The unbiased estimator for this correlation is the Landy–Szalay estimator (Landy & Szalay, 1993) shown in Equation 1.
@@
Our paper executes three experiments that are designed to explore three themes: model uncertainty, differentiability through an estimator, and differentiability though a surrogate.
@@
    Model Uncertainty and PSFs ($\S$5): For this experiment, we use ShOpt.jl (Berman et al., 2024; Berman & McCleary, 2024) to produce PSF models for the COSMOS-Web NIRCam images (Casey et al., 2023). Using our proposed algorithm, we then compute $\rho$ statistics. Finally, we assess the epistemic uncertainty associated with our model and compare it to the uncertainty we would get with traditional bootstrapping techniques. Our PSF modeling experiment is geared towards observers, who do not have access to differentiable forward models of PSF residuals.
@@
    Differentiability and Gravity Simulations ($\S$6): For this experiment, we forward model gravitational interactions using a series of ordinary differential equations. We then compute correlation functions on the resulting catalog and test different techniques for differentiating the correlation function outputs with respect to the cosmological parameters that determined the simulation. Our gravitational simulation experiment is designed with theorists in mind. Theorists can produce multiple realizations of their model through simulations that are derived from different cosmological parameters. Here, we assess the feasibility of differentiating a correlation function with respect to $G$.
@@
    Surrogates and Galaxy Intrinsic Alignment ($\S$7): For this experiment, we use an existing surrogate model to relate astrophysical model parameters to IA (Pandya et al., 2024, 2025). Using this surrogate, we use Hamiltonian Monte Carlo (HMC) to find posterior distributions over astrophysical model parameters that are likely to correspond to correlations of zero. Our IA experiment shows how to compensate for forward models and estimators that are not easily differentiated, and are again geared towards theorists who are able to realize many instances of their forward model through inference in order to train a surrogate. Our usage of HMC leverages the differentiability of neural networks to accelerate the sampling procedure in comparison to traditional algorithms like Metropolis–Hastings (Betancourt, 2017; Robert et al., 2004).
The remainder of the paper is organized as follows: $\S$2 discusses the relevant notation and theory for $\rho$ statistics and IA, $\S$3 follows with an outline of the data products used in our study, $\S$4 lays out additional notation as well as details for the algorithm used in this paper, and $\S$5 -7 presents the results of the experiments described above followed by our conclusions in $\S$8.
@@ 2.1 PSF Modeling
where $e_{\rm PSF}$ is the ellipticity of the real PSF, i.e., the star ellipticity, $T_{\rm PSF}$ is the size of the real PSF, $\delta e_{\rm PSF}$ is the difference between the ellipticity of the real and model PSFs at position $\bm{x}$, and $\delta T_{\rm PSF}$ is the difference between the sizes of the real and model PSFs at position $\bm{x}$. Brackets denote averages over all pairs within a separation $\bm{\theta}$, and asterisks denote complex conjugates. $T_{\rm PSF}$ is nominally defined as $2\sigma^{2}$, where $\sigma^{2}$ is the variance of the best fit elliptical Gaussian to the PSF.
@@
where $w_{i/j}$ represents the weight uncertainties of object $(i/j)$, $\epsilon_{t,i/j}$ represents tangential shear of object $(i/j)$ and $\epsilon_{\times,i/j}$ represents the cross shear of object $(i/j)$ (Kilbinger, 2015; Schneider et al., 2002). The complex components of the products in Equations 2 - 6 are known to average out to zero due to parity symmetry (Rowe, 2010; Schneider et al., 2006), and are thus ignored in the estimator.
@@ 2.2 Galaxy Intrinsic Alignment
Intrinsic alignment describe correlations between galaxy orientations and the underlying distribution of dark matter where they are embedded; see Lamman et al. (2023) for a review. The three relevant IA correlation functions are the galaxy position-position $\xi(r)$, position-orientation $\omega(r)$, and orientation-orientation $\eta(r)$ correlation functions, and are estimated according to Equations 9-11.
@@ 3.1 Point Source Catalog
The point sources used for generating our PSF models come from the SE++++ (Bertin & Arnouts, 1996) COSMOS-Web catalog (Casey et al., 2023, Shuntov et al. in prep). We restrict ourselves to imaging from the F115115W NIRCam (Rieke et al., 2003, 2005; Beichman et al., 2012) wavelength filter, as the PSFs in this wavelength are the closest to being well-approximated by a Gaussian (Berman et al., 2024). $\rho$ statistics and the related Hirata–Seljak–Mandelbaum (HSM) adaptive moments (Hirata & Seljak, 2003; Mandelbaum et al., 2005) are based off of the Gaussian approximation of the PSF, and so the use of $\rho$ statistics in our experiments is most appropriate in this wavelength. PSF models are obtained using ShOpt.jl (Berman et al., 2024; Berman & McCleary, 2024). Our choice of PSF fitter is discussed in more detail in Appendix A. Our ShOpt fits can be seen in Figure 3. The ellipticites of the objects in our data set are also visualized in Figure 4. We use the same configuration file as found in the appendix of Berman et al. (2024). Our data set statistics for PSF modeling are summarized in Table 1.
@@ 3.2 Simulated Galaxy Catalog
There are plenty of widely-used physically realistic N-body simulations, using publically available codes such as Gadget-2/3 (Springel, 2005), ART (Kravtsov et al., 1997), RAMSES (Teyssier, 2002), and Enzo (Brummel-Smith et al., 2019; Bryan et al., 2014), see Vogelsberger et al. (2020) for a review. However, this work aims only to have a simple differentiable forward model. Instead of using Gadget-2 or similar tools, we write our simulations from scratch in Julia and Python and time evolve objects according to ordinary differential equations from Newtonian mechanics. Our main implementation uses Jax (Bradbury et al., 2018) with Diffrax (Kidger, 2021). We specify starting positions and velocities for objects $\vec{x},\vec{v}]$, which are then evolved via $\frac{d\vec{x}}{dt}=\vec{v}$, $\frac{d\vec{v}}{dt}=\vec{a}$, and $\vec{a}=-\frac{GM}{|\vec{x}|^{3}}\vec{x}$. We begin by randomly initializing galaxies with some mass $M$, velocity $\vec{v}$, and initial position $\vec{r}$. These galaxies exist in a gravitational field where the galaxy-galaxy gravitational attraction is assumed to be negligible compared to the attraction toward the center of the field. Thus, we can solve each ODE system in Diffrax independently and use the vmap function to efficiently time evolve each object. While our toy simulation is not very physically meaningful, this approach outputs catalogs that are differentiable with respect to the input cosmology, which in this case is just the gravitational constant $G$. The number of galaxies in the simulation can vary, allowing us to study the efficacy of our algorithm as a function of catalog sizes.
@@ 3.3 Intrinsic Alignment Catalog
$\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt.$ (14)
The contributions of Van Alfen et al. (2024) introduce a model of IA within a HOD-based framework. Specifically, the IA implementation includes a two-parameter family governing central and satellite alignment strengths ($\mu_{\rm cen}$ and $\mu_{\rm sat}$). The work explores different forms of alignment strength (e.g., constant, radially dependent), and demonstrates that this parameterization can effectively model IA in both dark-matter only simulations which include galaxies via HOD modeling, and within the TNG300 suite of hydrodynamic simulations.
For each of the seven input parameters, ten realizations of the model were produced. Once the catalogs were generated, the three IA correlation functions were estimated according to Equations 9 - 11. This led to a data matrix of the shape $(110,526;7;100;3;20)$. In other words, we have $110,526$ tuples of seven input values that produced ten realizations of three correlation functions that each occupy twenty radial bins. This catalog is used to train the IA emulator (IAEmu) surrogate model introduced in Pandya et al. (2024). IAEmu is a neural network based emulator designed to predict galaxy IA statistics from a given HOD, with the alignments parameterized according to the two parameter family $\mu_{cen}$ and $\mu_{sat}$ (Van Alfen et al., 2024). These $110,526$ tuples are broken down in training, validation, and testing sets for IAEmu. This is summarized in Table 2.
@@ 4 Algorithm Overview
In this section, we outline our algorithms used to estimate correlations. Our algorithm has three variants, one geared toward estimating clustering uncertainties and the other two towards differentiability. Variations on the algorithm geared toward estimating clustering uncertainties are needed because estimating uncertainty requires sampling, which is incompatible with differentiation. There are a variety of ways to adapt sampling to be a differentiable process, which we explore in depth in §6. Throughout this work, we use most of the formalism laid out in Bezdek et al. (1984). Our formalism differs in that it uses a distance function to compute angular distances on the sky rather than using a matrix to define a norm. We also define the quantity matrix, which allows us incorporate the quantities we wish to correlate into the original formalism. For details on the fuzzy sets formalism, see Zadeh (1965) and (page 38, Aluffi, 2021). Algorithm 1 first uses K-means++++ (Kapoor & Singhal, 2017) to initialize a series of cluster centers. This algorithm adds initial cluster centers sequentially by considering the distances to centers that have already been determined. This initialization has been shown to help clustering algorithms and is better than choosing initial cluster centers at random (Kapoor & Singhal, 2017). We next use the fuzzy-c-means algorithm (Bezdek et al., 1984) to obtain final cluster centers and weight assignments that minimize the functional
@@
While this does not have a compact matrix representation, it is stable at all numerical scales (Vincenty, 1975). In this way, we are sacrificing speed for stability. This design choice is discussed more in $\S$6. The two argument $\arctan$ function is used to constrain the solution to the correct quadrant. Our usage of the Vincenty formula also assures that our distance estimates are consistent with Astropy (Robitaille et al., 2013), which also uses the Vincenty formula for this task.
@@
    Initialize centers via K-means++++ and weights via random initialization.
@@
1: Initialize centers $v$ using K-Means++++
@@
Figure 7: An example weighting for estimates for a distance bin bounded below by 11 and above by 22 for various sharpnesses.
@@
1: Initialize centers $v$ using K-Means++++
@@
As we will see, there are some computational hurdles that any differentiable estimator will inevitably face that can make differentiability intractable in practice. These can be circumnavigated using a surrogate model, as outlined in Algorithm 3.
@@
5 Model Uncertainty
In this section, we discuss the results of Algorithm 1 on our point source catalog dataset, which we compare with a bootstrapping baseline to obtain epistemic and aleatoric uncertainties.
6 Differentiable Forward Model
For both approaches b and c (as defined in $\S$4), the naive approach to computing the gradient $\frac{\partial\xi}{\partial\alpha_{i}}$ or $\frac{\partial\xi}{\partial\delta_{i}}$ would involve tracking the computation graph through the entire fuzzy-c-means algorithm, which can sometimes take over $100$ iterations. Another complication is that the distance calculations and normalization computations cannot be expressed through matrix multiplication, making it difficult to accelerate the gradient calculations with a GPU. For approach $b$, taking the gradients in this manner proved to be prohibitively expensive via reverse mode automatic differentiation, for the reasons outlined above. With forward mode differentiation, the gradient calculation was on the order of a minute. This is expected, since we are differentiating several output bins with respect to $1$ input (gravity).
7 Differentiable Surrogates
In this section, we show how Algorithm 3 can be used to differentiate through a correlation function and present a compelling science case that supports its utility. In particular, we utilize the surrogate model developed in Pandya et al. (2024, 2025), IAEmu, which estimates galaxy IA correlations from an underlying set of HOD parameters. We use IAEmu to optimize model parameters that minimize the IA correlations, to study regions of parameter space that correspond to minimal IA contamination and are thus of physical interest. IAEmu serves as an emulator for Halotools-IA, which extends traditional HOD modeling to include IA information through a two parameter family, $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$, which govern central and satellite alignment strengths, respectively. More specifically, IAEmu outputs the galaxy position-position ($\xi$), position-orientation ($\omega$), and orientation-orientation ($\eta$) correlations as well as estimates of the aleatoric uncertainty (shape noise) of each correlation. This noise is not physical, and corresponds to variance across realizations of the underlying HOD due to limited cosmological volumes when conducting the HOD simulations.
We pose this optimization problem because IA often act as a contaminant for weak lensing measurements, and moreover, there are known regions of parameter space that lead to low correlations. Thus, with this experiment, we are able to demonstrate that a differentiable forward model allows one to learn something about the underlying physics of IA. In other words, we are able to recover a known parameter space even without hard-coding the physical laws into the surrogate directly. This is distinct from Pandya et al. (2025), who use the emulator to find the IA parameters that best explain a set of observations rather than the parameters that result in the minimized correlation function.  [^3] The exact correlation that would be subtracted is the gravitationally lensed shape - intrinsic shape $(GI)$ correlation (Lamman et al., 2023); however, the strength of that correlation is closely related to that of $\omega$.
To isolate regions of parameter space in $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$ that minimize the IA correlation $\omega(r)$, we utilize Hamiltonian Monte Carlo (HMC). HMC is a gradient based Monte Carlo technique which leverages the differentiability of NNs and uses gradient information to efficiently traverse the input parameter space towards regions that maximize the posterior probability. For a general review of HMC, see Betancourt (2017). Our HMC experiment mirrors Pandya et al. (2025). We run our HMC with a No U-Turn Sampler (NUTS) (Hoffman et al., 2014) with $1000$ warmup steps and $1000$ samples. We fix the HOD parameters according to the fiducial values present in IllustrisTNG300-1 (Nelson et al., 2019) and Table C1 of Van Alfen et al. (2024). This gives us three sets of HOD parameters, which each correspond to a different stellar mass cutoff in galaxies. We impose uniform priors of [1,1][-1,1] on $\mu_{\rm cen}$ and $\mu_{\rm sat}$ and use a multivariate normal distribution centered at zero as our likelihood, with a covariance given by the IAEmu aleatoric uncertainty prediction. In this way, we encourage HMC to arrive at regions of parameter space that minimize the $\omega$ correlation, while also incorporating the inherent shape noise into the likelihood.